An experiment to measure the enthalpy change for the reaction of aqueous copper(II) sulfate, CuSO4(aq) and zinc, Zn(s) was carried out in a coffee cup calorimeter as follows.
Cu2+(aq) + Zn(s) → Cu(s) + Zn2+(aq)
50.0 cm3 of 1.00 mol dm–3 copper(II) sulfate solution was placed in a polystyrene cup and its average temperature after three readings was noted to be 25OC and the zinc powder was added . The final temperature was noted to be
91.5 OC. Calulate the heat of the reaction in kJ
1
Expert's answer
2021-09-06T02:35:47-0400

In your case, you know that the total heat capacity of the system, which includes that of the solution, is

500 J K−1

This tells you that in order to increase the temperature of the system by

1 K

, you need to provide it with

500 J

worth of heat.


Since the temperature of the solution increased by

66.5 K

, the amount of heat absorbed must have been


66.5 × 500/1 = 33250J

qsys=−n⋅ΔH

, where


n - the number of moles of copper sulfate that take part in the reaction.

ΔH - the enthalpy change of reaction per mole.


Use the solution's volume and molarity to determine how many moles of copper sulfate were present

1 × 50 × 10-3L = 0.05 mol

∆H = 33250/0.05

= 665000

= 665kJmol-1